The wave motion hits the interface of the two substances at an inlet angle of 36.0 ° and bends at an angle of refraction of 61.9 °. How does it behave when it hits the same interface at an entrance angle of 53.0 °?


1
Expert's answer
2021-05-02T18:32:04-0400
"n=\frac{\sin{61.9}}{\sin{36}}=1.5\\\theta=90\degree"